If a simple pendulum of length l oscillates through an angle α on either side of the mean position then find the angular velocity dt/dθ
of the pendulum where θ is the angle which the string makes with the vertical.
1
Expert's answer
2020-03-03T14:57:43-0500

The equation of motion for the pendulum:

"\frac{d^2\theta}{dt^2}=-\frac{g}{l}sin\theta, \;\;\theta (0)=\alpha,\;\;\frac{d\theta (0)}{dt}=0."

For small angles "\theta" we can use the approximation "sin\theta=\theta."

Thus, the solution of the initial value problem

"\frac{d^2\theta}{dt}=-\frac{g}{l}\theta,\;\; \theta (0)=\alpha,\;\;\frac{d\theta}{dt}=0"

is "\theta=\alpha cos(\sqrt{\frac{g}{l}}t)."

"v(t)=\frac{d\theta}{dt}=-\alpha\sqrt{\frac{g}{l}}sin(\sqrt{\frac{g}{l}}t)".